NASA ISS On-Orbit Status 2 December 2014

©NASA

NASA ISS On-Orbit Status 2 December 2014.

Today: EXPRESS Rack (ER) Gateway Configuration: Virts configured the ER6, 7, 8 laptops to communicate with the Payload Ethernet Hub/Gateway (PEHG)-2 allowing remote desktop access from the ground to the laptops via the ISS Ku-Band link.

3D Printing in Zero G: Cristoforetti removed and stowed a 3D printed coupon that was created yesterday. Afterwards, ground teams commenced printing the next coupon. The 3D printing in Zero-G technology demonstration experiment aims to validate extrusion-based additive manufacturing technology's ability to function in a micro-gravity environment.

Node 3 Carbon Dioxide Removal Assembly (CDRA) Bed Removal: Wilmore and Virts performed the complex task to remove the Node 3 CDRA from its rack and extracted the Desiccant/Sorbent Beds 201 and 202 from the unit. Cleaning of the beds is scheduled to occur Wednesday and Thursday of this week. Pressures across the rear and front desiccant/adsorbent beds of the Node 3 CDRA have been increasing since the previous cleaning in June of this year. The reduction in flow is due to flaking adsorbent material (Zeolite) inside the beds. The main goal of this week's task is to separate each desiccant/adsorbent bed into its two parts, clean the Zeolite, change out a filter, and then re-assemble the beds. Lab CDRA is operating this week to maintain CO2 within limits.

Service Module Voltage/Current Stabilizer [CHT] Testing: Cristoforetti setup two Station Support Computers (SSCs) to measure current in CHTs. During setup, three current clamps were configured with new batteries then connected to USB Oscilloscopes and had their zero set point verified. Samokutyaev assisted with power connection for SSC Laptops and identification of wires to be tested. The goal of CHT testing is to assess their performance by capturing high speed 2 channel oscilloscope electrical current data at the input (ISS Primary Power 151-160VDC) and output (28VDC) of each CHT simultaneously over a 4 day period, while forcing them into different modes, "Current Regulation" and "Voltage Regulation".

Today's Planned Activities
All activities were completed unless otherwise noted.

Hematocrit Test
HRF Urine Sample Collection & Insertion Into MELFI
Regeneration of БМП Ф2 Micropurification Cartridge (start)
Opening СНТ pit and Med locker
Verification of anti-virus scan results on Auxiliary Computer System (ВКС) laptops
24-hour ECG (start)
SM Power Supply System Feeder Test: Connecting SSC Laptop to station power outlet.
IFM - Assembly of Measurement configuration using 1- and 2-channel USB oscilloscope, hardware setting adjustments
Removal of two КЛ-152М TV cameras and light units from ТК 715 СА
SM Electric Power Supply System Feeder Test: Connecting current probes (part A and part B)
SM CO measurement using CMS unit
HRF - Hardware Setup
RGN-REC TK - Transfer and Replacement
ARED Guide Rail Inspection
EHS - TEPC Removal
MRM1 Fan Screen Cleaning (Group B)
SM Electric Power Supply System Feeder Test: Switching current probe (part B)
SM Comm Subsystem [СТТС] routine maintenance
WRS - Recycle Tank Fill from EDV
IMS Delta File Prep
Node 3 Carbon Dioxide Removal Assembly (CDRA) Removal for Cleaning and Replacement
PAO Event
SM Electric Power Supply System Feeder Test: Switching current probe (part B)
WRS - Recycle Tank Fill from EDV
MELF1 - Crew inserts Ice Bricks
СТТС Configuration for MRM1 comm
CDRA - BED Removal
SM Electric Power Supply System Feeder Test: Switching current probes (part A and part B)
WRS - Recycle Tank Fill from EDV
KHIMIYA OBRAZOVANIYE. Sfera-S Session
VHF2d test from SM (using primary СТТС sets )
RGN REC-TNK - Removal of depress hose for nominal operations
СТТС comm assets switchover to a backup set
Filling (Separation) of EDV (KOV) Lid No.10 п.36 (ФГБ1ПГО_4_404_1) from EDV lid 1104 (ФГБ1ПГО_1_109) for Elektron system
WRM Condensate Pumping
WRS - Recycle Tank Fill from EDV
VHF2d Test from SM (on backup comm sets)
3DP - Printed Coupon Removal and Stowage
Closing СНТ pit
БМП Ф2 Absorption Cartridge Regeneration (termination)
Completed Task List Items

ARED Detent Flip
PIP Lab Server Relocation
Ground Activities

All activities were completed unless otherwise noted.

JEM RMS SFA stow and deactivation
Relocate SPDM from MBS to the Lab PDGF [In Work]
Three-Day Look Ahead:

Wednesday, 12/03: CDRA Bed cleaning and Insulation Replacement, Biolab TCU Cleaning and Silica Bag Exchange
Thursday, 12/04: CDRA Bed Cleaning part 2 and reassembly, 3D Printing, SKIN B, Robotics OBT
Friday, 12/05: 3D Printing, CDRA install and leak check, Glacier 2 Desiccant Pack Swap
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - On
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Operate
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Standby
Major Constituent Analyzer (MCA) Lab - Stop
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Normal
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.